A solution of ephedrine has a pH of 11.8. What is its initial concentration and the percent ionization? Kb=1.4x10^-4. Ephedrine = C10H15ON
Answer #1

Kb = 1.4 pk saь pH = 11-8 polH_ 2.2 c=2 poh = + (pks loge) 2-2 = 4 (3-85-logc) 4:4=3.85-loge = -log (1-4X10%) =4-log(1-4) = 3